INTRINSIC MOLECULAR IONIC CENTERS IN RARE GAS SOLIDS

\underline{A. N. OGURTSOV}, E. V. SAVCHENKO, O. N. GRIGORASHCHENKO, O. M. SOKOLOV, Verkin Institute for Low temperature Physics & Engineering, 47 Lenin Avenue, 310164 Kharkov, Ukraine.

Electronically induced formation of intrinsic ionic centers in the configuration of rare gas molecular ions (R_{2}^{+}) was studied by low temperature spectroscopy methods. Based on the cathodoluminescence investigation of the ionic excimers (R_{2}^{2+}) ^{a}, thermoluminescence data on self-trapped holes (R_{2}^{+}) ^{b} and selective excitation by synchrotron radiation of states involved ^{c}, the set of lattice reactions were analyzed: 1) free exciton ionization accompanied by trapping of charge carriers: R_{ex}^{∗} → R_{2}^{+} + e^{−}; 2) ‘dressing’ of trapped hole by exciton with subsequent H-band emission: R_{2}^{+} + R_{ex}^{∗} → R_{2}^{+∗} → R_{2}^{∗} + hν(H-band); 3) bulk recombination of trapped hole with electron following by exciton self-trapping and M-band emission: ^{bulk}R_{2}^{+} + e^{−} → ^{bulk}R_{2}^{∗} → R + R + hν(M-band); 4) surface recombination of trapped hole with electron following by desorption of excimers and W -band emission: surf R_{2}^{+} + e^{−} → ^{desorb}R_{2}^{∗} → R + R + hν(W -band).

aO. N. Grigorashchenko, S. A. Gubin, A. N. Ogurtsov, E. V. Savchenko, J. Electron Spectrosc. Relat. Phenom. 79, 107 (1996).

bE. V. Savchenko, A. N. Ogurtsov, S. A. Gubin, O. N. Grigorashchenko, In: Proc. 2nd Int. Conf. on Excitonic Processes in Condensed Matter (EXCON’96) (Dresden University Press, Dresden 1996) p.207-210.

cJ. Becker, A. N. Ogurtsov, M. Runne, E. V. Savchenko, G. Zimmerer, HASYLAB Annual Report 1996 (DESY, Hamburg 1997)

in press.